# EvaNet: Elevation-Guided Flood Extent Mapping on Earth Imagery

### v2

@@ 1 Introduction
Climate change is drastically increasing the intensity and occurrence of floods [?]. In just the last two decades, flooding has negatively impacted over 2.3 billion people, often disproportionately affecting vulnerable communities  [?]. Therefore, accurate and timely mapping of flood extent can be crucial for effectively planning rescue and rehabilitation efforts  [?]. However, sending field crews on site for inspection is prohibitively slow, while obtaining data from a limited number of groundwater sensors or UAVs cannot provide a global coverage. Therefore, we target the accurate and timely mapping of flood extent from high-resolution satellite imagery that provides a global coverage.
There has been a lot of research on using remote-sensing satellite imagery for flood extent mapping. Two commonly used satellite types are radar (active sensor) and optical (passive sensor). For example, Sentinel-1 [?] and Sentinel-2 [?] are radar and optical satellites, respectively, that produce high resolution earth imagery at high revisit times [?]. Radar is more resilient to noises like clouds, but traditional approach to generating flood maps from radar imagery relies on semi-automated tools, [?] or thresholding techniques [?] that are slow and lack generalizability, often requiring expert intervention for parameter tuning. In this paper, we focus on the optical satellite imagery, since radar data such as that from Sentinel-1 [?] can also be converted into RGB channels [?]. Moreover, both [?] and [?] found that using Sentinel-2 data only yields the best segmentation performance for deep learning models, as compared with when Sentinel-1 data is involved.
@@
Thanks to the availability of digital elevation model (DEM) data from sources such as USGS, existing deep learning approaches use the elevation map to improve segmentation performance by treating the map simply as an additional input channel of U-Net [?] just like a regular spectral channel [???]. The physical law of gravity inherent to elevations as illustrated in Figure 1 (see its caption) is not encoded into the model design.
Encoding physical law in model design can help better recover the true states (“flooded” or “dry”) of noisy pixels covered by clouds and tree canopies, which are often ambiguous just based on spectral features, and it has shown success in graphical models for flood mapping [????]. We propose to encode the physics-aware constraints into the more powerful deep learning models to achieve better performance.
@@
  • We extensively explored various approaches to integrating the elevation map as an input to the convolution operations on the encoder and decoder sides of U-Net, and find that using an independent convolution branch over the elevation map (called the elevation path) to regulate the flow of spectral features by the gated linear unit (GLU) [?] gives the best performance by striking a balance between model simplicity (to prevent overfitting) and model expressiveness.
@@ 2 Related Work
Image Segmentation. Flood extent mapping is a semantic segmentation task, which classifies each pixel in an image into a semantic class (flood or dry in our problem). Besides semantic segmentation, there are other segmentation tasks such as instance segmentation [?] which finds the object pixels in the bounding box of an object detected by an object detection model, and panoptic segmentation [?] which combines semantic segmentation and instance segmentation.
@@
Flood Extent Mapping by Deep Learning. A number of deep learning models for flood extent mapping have been explored by the remote sensing community recently. For example, [?] first applies FCN [?] on only the RGB data to obtain an initial flood map (Stage 1); then the work uses water level data from USGS combined with the DEM data to detect water pixels under tree canopy, by using a method called region growing (Stage 2); the result from Stage 2 is then used to improve the FCN-based flood extent for vegetated areas (Stage 3). We remark that elevation data was not used in Stage 1, and our EvaNet can serve as a drop-in replacement of FCN to improve the performance of Stage 1.
@@
As another example, [?] uses spectral bands SWIR2, NIR and red values from Sentinel-2 to compute new indices as U-Net input, and DEM data is treated simply as an additional channel. Also, [?] focuses on the application of onboard segmentation with small, nano satellites to reduce revisit time to disaster areas; they input 13-band Sentinel-2 image into U-Net for segmentation without using DEM data. While these works from the remote sensing community include new input modes that can extend our work, our EvaNet targets the more fundamental problem of effectively integrating DEM data with spectral features guided by the physical law of gravity, and can serve as a drop-in replacement of their naïve U-Net or FCN models to improve performance.
@@
Non-Deep-Learning Methods. Hidden Markov tree (HMT) [?] is a state-of-the-art graphical model for flood extent mapping. HMT captures the directed spatial dependency based on flow directions across all locations by a reverse tree structure in the hidden class layer, which is similar to the hidden state sequence in hidden Markov models (HMM); spectral features are read out from the class state (dry or flooded) based on a learned multivariate Gaussian distribution. Later, hidden Markov contour tree (HMCT) [?] further captures complex contour structures on a 3D surface.
@@
Both HMT and HMCT are for transductive learning where training samples are taken near the test region, or where a complete elevation map plus only limited spectral features in a target region are observed [?]. More recently, inductive learning with HMT is explored with the help of U-Net. [?] proposes to use the inference stage of HMCT only as a post-processor of U-Net to cleanse its predicted noisy posterior class probability at all the pixels. The resulting model is called HMCT-PP. Our EvaNet can also serve as a drop-in replacement of the U-Net in HMCT-PP.
@@
Knowledge-Guided Machine Learning (KGML). Unlike our problem that focuses on the Earth imagery at one snapshot to quickly map out flood-inundated area for disaster response, KGML research is usually formulated as a forecast problem fitting historical data or fitting a simulation model (e.g., based on PDE), so requires many temporal snapshots. The goal of KGML is more on geosciences to study climate and environmental phenomena/process for future disaster warning and mitigation, often using meteorological factors such as air temperature, precipitation, and wind speed. We discuss the related works on KGML in the appendix.
@@ 3.1 Elevation-Guided Loss
Note that a pixel may be unlabeled since it is covered by tree canopy. We provide domain experts with a 3D visualization tool where they can navigate the terrain surface by shifting and rotation, and where they can annotate the pixels using tools like ‘brush’ if they are confident of the pixel labels. Elevation-guided breadth-first search (BFS) [?] was adopted to speed up annotation, where (i) when an annotator marks an individual pixel $\mathbf{p}$ as flooded, the label propagates to nearby pixels with lower elevations by ‘pit-filling’ breadth-first search stopping when reaching pixels with elevation higher than $h(\mathbf{p})$, and (ii) when an annotator marks an individual pixel $\mathbf{p}$ as dry, the label propagates to nearby pixels by ‘hill-climbing’ breadth-first search stopping when reaching pixels with elevation starting to drop. The labels of some pixels covered by tree canopy may be derived in this way from nearby pixels using the physical law of gravity.
@@
Figure 5 in our appendix shows the pixel distribution among the 5 cases on our datasets, where Cases 2.1 and 3.1 are dominating where the physical constraints apply, so we expect sufficient regularization.
@@
However, we find that these weighting schemes are not better than binary weighting. The results of comparison can be found in Table 7 of the appendix.
@@
where hyperparameter $\lambda$ balances the importance between $\mathcal{L}_{CE}$ and $\mathcal{L}_{eva}$. U-Net simply uses $\mathcal{L}=\mathcal{L}_{CE}$ (i.e., $\lambda=0$). By setting $\lambda=1$, we obtain a representative hybrid loss $\mathcal{L}=\mathcal{L}_{CE}+\mathcal{L}_{eva}$. Another extreme is using $\mathcal{L}=\mathcal{L}_{eva}$ (i.e., $\lambda=\infty$), which could be a bit surprising since we exclude $\mathcal{L}_{CE}$ and only use the regularization term to penalize the violation of physical constraints. Surprisingly, this last loss turns out to perform well in our experiments (c.f., Table 6 in our appendix), probably because Cases 2.1 and 3.1 are well spread out on the entire terrain (c.f., Figure 5 in our appendix). In other words, $\mathcal{L}_{eva}$ by itself serves as a good loss function for flood mapping.
@@ 3.2 Elevation-Regulated Convolutions
As Figure 4 shows, at the encoder side, $\mathbf{Y}$ is downsampled using $2\times 2$ max pooling which is commonly used for spectral tensors, and $\mathbf{Y}_{e}$ is downsampled using $2\times 2$ average pooling as a lower-resolution terrain approximation. We tried other pooling combinations but this setting gives the best predictions as demonstrated in Table 8 of the appendix. At the decoder side, we use $3\times 3$ transpose convolution for both $\mathbf{Y}$ and $\mathbf{Y}_{e}$.
@@ 4 Experiments & Results
Data. We obtain high-resolution aerial imagery from NOAA National Geodetic Survey during Hurricane Matthew in North Carolina (NC) in 2016 [?]. The accompanied DEM data are obtained from the University of North Carolina Libraries [?]. Futhermore, to test if our models trained on the Matthew 2016 dataset can generalize to a different flooding event, we also obtained two more regions in Baytown, TX, and Richmond, TX during Hurricane Harvey in 2017 from [?], and the corresponding DEM data was obtained from USGS data downloader [?]. The aerial imagery from both events are 0.3 m $\times$ 0.3 m in resolution, the DEM data for Matthew 2016 are 6 m $\times$ 6 m and for Harvey 2017 are 3 m $\times$ 3 m in resolution. We resampled them into a resolution of 2 m $\times$ 2 m, which is fine enough for the purpose of flood extent mapping. Table 2 shows the regions we used, where R1–R5 are from Matthew 2016, and R6–R7 is from Harvey 2017. While high-resolution Earth imagery and DEM data are abundant, the ground-truth flood maps need to be annotated by domain experts. So, we provide a 3D visualization tool to domain experts for annotating pixels.
In our appendix, Figure 6 shows the imagery for Kinston, NC during Hurricane Matthew in 2016, its annotation where pixels are labeled as flooded or dry, and the normal-time imagery from Google Earth. We expect that inputting the normal-time image into a model is effective in helping identify flooded regions by providing a contrast. This is confirmed our experiments reported in Table 10 of our appendix, which also aligns with findings from prior works [??]. Some pixels are ambiguous (e.g., covered with tree canopies) so remain unlabeled. Unlabeled pixels are not included in the loss computation during training, and not included during test when calculating the evaluation measures. As Table 2 shows, most pixels of each region are annotated. Without loss of generalization, in Table 2, we use R1 and R2 for training, and use R3, R4, R5, R6 and R7 for test.
@@
For physics-guided graphical models, we compare with hidden Markov tree (HMT) [??]. Since HMT uses transductive learning, we use the same region for training and test. For U-Net and EvaNet, we use two regions R1 and R2 in Table 2 for training and the other regions for test. For example, when reporting the performance measures on R3, HMT is trained on R3 and tested on R3, while all other models are trained with the patches of R1 and R2, and then tested on patches of R3 and stitched.
By default, EvaNet uses only $\mathcal{L}_{eva}$ in Eq (2) as the loss function, where $w(\mathbf{p},\mathbf{p}_{n})$ uses the binary weighting scheme in Eq (4). We use 3 convolution blocks of Eva-Conv for both the encoder and the decoder as shown in Figure 4, where the input has 7 channels (disaster-time and normal-time RGB plus elevation map), and the encoder downsamples the spectral path using max pooling, and downsamples the elevation path using average pooling. These default configurations are found to be the best-performing, as demonstrated by our experiments in the appendix (see Tables 610). We trained our EvaNet models and the U-Net baselines on a cluster with NVIDIA-P100 GPUs for 100 epochs with a learning rate of $1\times 10^{-7}$ and batch size of 4. These default settings are found to be the best-performing through extensive tests.
Performance Measures. Accuracy, precision, recall and F1-score are commonly used measures for binary classification. Following prior works such as [??], we consider our flood segmentation problem as a pixel-wise binary classification problem, and report the above four measures in two contexts: (1) “Dry” where dry (resp. flood) is the positive (resp. negative) class, and (2) “Flood” where dry (resp. flood) is the positive (resp. negative) class. These measures are computed only from labeled pixels in the test images.
Comparison with Baselines. Here, we consider 3 representative baselines: (1) HMT [?], which is a graphical model for transductive learning that considers both the disaster-time RGB image and the elevation map, (2) U-Net 3C, which is a U-Net model that only considers the disaster-time RGB image, (3) U-Net 7C, which is a U-Net model that considers both the RGB images and the elevation map. They are compared with EvaNet 7C in Table 3. We skip 4C models since they are dominated by 7C ones (c.f. Table 10 in our appendix).
@@
Ablation Study. To verify the effectiveness of both our Eva-Conv and Eva-Reg designs, we conduct an ablation study using 7-channel input, by exploring the performance of 4 models: (1) U-Net 7C (i.e., with loss $\mathcal{L}_{CE}$), (2) U-Net with our loss function $\mathcal{L}_{eva}$ (denoted by U-Net 7C + Eva-Reg), (3) U-Net 7C with our new Eva-Conv operations in both encoder and decoder (denoted by U-Net 7C + Eva-Conv), and (4) EvaNet 7C, i.e., with both Eva-Conv and Eva-Reg. Without loss of generality, Table 4 shows the performance results on test region R4, where we can see that (i) both Eva-Reg and Eva-Conv improve the model accuracy and F1-score as compared with U-Net 7C, (ii) Eva-Reg (i.e., our new loss $\mathcal{L}_{eva}$) improves the performance more significantly, and (iii) our EvaNet 7C model combining both designs provides the highest accuracy and F1-score. In our appendix, we show that our techniques are effective also in other ConvNet architectures using Table 11, and we also visually inspect the flood maps from different models in Figure 7 to confirm our conclusions.
Effect as a U-Net Drop-in Replacement. We also find that EvaNet can improve the performance of an existing model such as HMCT-PP [?] (i.e., U-Net + HMCT) by replacing its U-Net component. Due to space limit, please refer to Table 12 in our appendix for more details.

### v3

@@ 1 Introduction
Climate change is drastically increasing the intensity and occurrence of floods Matgen et al. (2020). In just the last two decades, flooding has negatively impacted over 2.3 billion people, often disproportionately affecting vulnerable communities  Wahlstrom et al. (2015). Therefore, accurate and timely mapping of flood extent can be crucial for effectively planning rescue and rehabilitation efforts  Oddo and Bolten (2019). However, sending field crews on site for inspection is prohibitively slow, while obtaining data from a limited number of groundwater sensors or UAVs cannot provide a global coverage. Therefore, we target the accurate and timely mapping of flood extent from high-resolution satellite imagery that provides a global coverage.
There has been a lot of research on using remote-sensing satellite imagery for flood extent mapping. Two commonly used satellite types are radar (active sensor) and optical (passive sensor). For example, Sentinel-1 Copernicus (2014) and Sentinel-2 Copernicus (2015) are radar and optical satellites, respectively, that produce high resolution earth imagery at high revisit times Matgen et al. (2020). Radar is more resilient to noises like clouds, but traditional approach to generating flood maps from radar imagery relies on semi-automated tools, Rudner et al. (2019) or thresholding techniques Fraccaro et al. (2022) that are slow and lack generalizability, often requiring expert intervention for parameter tuning. In this paper, we focus on the optical satellite imagery, since radar data such as that from Sentinel-1 Copernicus (2014) can also be converted into RGB channels Fraccaro et al. (2022). Moreover, both Konapala et al. (2021) and Bonafilia et al. (2020) found that using Sentinel-2 data only yields the best segmentation performance for deep learning models, as compared with when Sentinel-1 data is involved.
@@
Thanks to the availability of digital elevation model (DEM) data from sources such as USGS, existing deep learning approaches use the elevation map to improve segmentation performance by treating the map simply as an additional input channel of U-Net Ronneberger et al. (2015) just like a regular spectral channel Beni and Gebrehiwot (2021); Konapala et al. (2021); Fraccaro et al. (2022). The physical law of gravity inherent to elevations as illustrated in Figure 1 (see its caption) is not encoded into the model design.
Encoding physical law in model design can help better recover the true states (“flooded” or “dry”) of noisy pixels covered by clouds and tree canopies, which are often ambiguous just based on spectral features, and it has shown success in graphical models for flood mapping Jiang and Sainju (2019); Sainju et al. (2022, 2020); Xie et al. (2018). We propose to encode the physics-aware constraints into the more powerful deep learning models to achieve better performance.
@@
  • We extensively explored various approaches to integrating the elevation map as an input to the convolution operations on the encoder and decoder sides of U-Net, and find that using an independent convolution branch over the elevation map (called the elevation path) to regulate the flow of spectral features by the gated linear unit (GLU) Dauphin et al. (2017) gives the best performance by striking a balance between model simplicity (to prevent overfitting) and model expressiveness.
@@ 2 Related Work
Image Segmentation. Flood extent mapping is a semantic segmentation task, which classifies each pixel in an image into a semantic class (flood or dry in our problem). Besides semantic segmentation, there are other segmentation tasks such as instance segmentation He et al. (2017) which finds the object pixels in the bounding box of an object detected by an object detection model, and panoptic segmentation Kirillov et al. (2019) which combines semantic segmentation and instance segmentation.
@@
Flood Extent Mapping by Deep Learning. A number of deep learning models for flood extent mapping have been explored by the remote sensing community recently. For example, Beni and Gebrehiwot (2021) first applies FCN Long et al. (2015a) on only the RGB data to obtain an initial flood map (Stage 1); then the work uses water level data from USGS combined with the DEM data to detect water pixels under tree canopy, by using a method called region growing (Stage 2); the result from Stage 2 is then used to improve the FCN-based flood extent for vegetated areas (Stage 3). We remark that elevation data was not used in Stage 1, and our EvaNet can serve as a drop-in replacement of FCN to improve the performance of Stage 1.
@@
As another example, Konapala et al. (2021) uses spectral bands SWIR2, NIR and red values from Sentinel-2 to compute new indices as U-Net input, and DEM data is treated simply as an additional channel. Also, Mateo-Garcia et al. (2021) focuses on the application of onboard segmentation with small, nano satellites to reduce revisit time to disaster areas; they input 13-band Sentinel-2 image into U-Net for segmentation without using DEM data. While these works from the remote sensing community include new input modes that can extend our work, our EvaNet targets the more fundamental problem of effectively integrating DEM data with spectral features guided by the physical law of gravity, and can serve as a drop-in replacement of their naïve U-Net or FCN models to improve performance.
@@
Non-Deep-Learning Methods. Hidden Markov tree (HMT) Xie et al. (2018) is a state-of-the-art graphical model for flood extent mapping. HMT captures the directed spatial dependency based on flow directions across all locations by a reverse tree structure in the hidden class layer, which is similar to the hidden state sequence in hidden Markov models (HMM); spectral features are read out from the class state (dry or flooded) based on a learned multivariate Gaussian distribution. Later, hidden Markov contour tree (HMCT) Jiang and Sainju (2019) further captures complex contour structures on a 3D surface.
@@
Both HMT and HMCT are for transductive learning where training samples are taken near the test region, or where a complete elevation map plus only limited spectral features in a target region are observed Sainju et al. (2020). More recently, inductive learning with HMT is explored with the help of U-Net. Sainju et al. (2022) proposes to use the inference stage of HMCT only as a post-processor of U-Net to cleanse its predicted noisy posterior class probability at all the pixels. The resulting model is called HMCT-PP. Our EvaNet can also serve as a drop-in replacement of the U-Net in HMCT-PP.
@@
Knowledge-Guided Machine Learning (KGML). Unlike our problem that focuses on the Earth imagery at one snapshot to quickly map out flood-inundated area for disaster response, KGML research is usually formulated as a forecast problem fitting historical data or fitting a simulation model (e.g., based on PDE), so requires many temporal snapshots. The goal of KGML is more on geosciences to study climate and environmental phenomena/process for future disaster warning and mitigation, often using meteorological factors such as air temperature, precipitation, and wind speed. We discuss the related works on KGML in our online appendix Sami et al. (2024).
@@ 3.1 Elevation-Guided Loss
Note that a pixel may be unlabeled since it is covered by tree canopy. We provide domain experts with a 3D visualization tool where they can navigate the terrain surface by shifting and rotation, and where they can annotate the pixels using tools like ‘brush’ if they are confident of the pixel labels. Elevation-guided breadth-first search (BFS) Adhikari et al. (2022) was adopted to speed up annotation, where (i) when an annotator marks an individual pixel $\mathbf{p}$ as flooded, the label propagates to nearby pixels with lower elevations by ‘pit-filling’ breadth-first search stopping when reaching pixels with elevation higher than $h(\mathbf{p})$, and (ii) when an annotator marks an individual pixel $\mathbf{p}$ as dry, the label propagates to nearby pixels by ‘hill-climbing’ breadth-first search stopping when reaching pixels with elevation starting to drop. The labels of some pixels covered by tree canopy may be derived in this way from nearby pixels using the physical law of gravity.
@@
Figure 5 in our online appendix Sami et al. (2024) shows the pixel distribution among the 5 cases on our datasets, where Cases 2.1 and 3.1 are dominating where the physical constraints apply, so we expect sufficient regularization.
@@
However, we find that these weighting schemes are not better than binary weighting. The results of comparison can be found in Table 7 of the appendix Sami et al. (2024).
@@
where hyperparameter $\lambda$ balances the importance between $\mathcal{L}_{CE}$ and $\mathcal{L}_{eva}$. U-Net simply uses $\mathcal{L}=\mathcal{L}_{CE}$ (i.e., $\lambda=0$). By setting $\lambda=1$, we obtain a representative hybrid loss $\mathcal{L}=\mathcal{L}_{CE}+\mathcal{L}_{eva}$. Another extreme is using $\mathcal{L}=\mathcal{L}_{eva}$ (i.e., $\lambda=\infty$), which could be a bit surprising since we exclude $\mathcal{L}_{CE}$ and only use the regularization term to penalize the violation of physical constraints. Surprisingly, this last loss turns out to perform well in our experiments (c.f., Table 6 in our appendix Sami et al. (2024)), probably because Cases 2.1 and 3.1 are well spread out on the entire terrain (c.f., Figure 5 in our appendix Sami et al. (2024)). In other words, $\mathcal{L}_{eva}$ by itself serves as a good loss function for flood mapping.
@@ 3.2 Elevation-Regulated Convolutions
As Figure 4 shows, at the encoder side, $\mathbf{Y}$ is downsampled using $2\times 2$ max pooling which is commonly used for spectral tensors, and $\mathbf{Y}_{e}$ is downsampled using $2\times 2$ average pooling as a lower-resolution terrain approximation. We tried other pooling combinations but this setting gives the best predictions as demonstrated in Table 8 of the online appendix Sami et al. (2024). At the decoder side, we use $3\times 3$ transpose convolution for both $\mathbf{Y}$ and $\mathbf{Y}_{e}$.
@@ 4 Experiments & Results
Data. We obtain high-resolution aerial imagery from NOAA National Geodetic Survey during Hurricane Matthew in North Carolina (NC) in 2016 NOAA (2016). The accompanied DEM data are obtained from the University of North Carolina Libraries NCSU (2023). Futhermore, to test if our models trained on the Matthew 2016 dataset can generalize to a different flooding event, we also obtained two more regions in Baytown, TX, and Richmond, TX during Hurricane Harvey in 2017 from NOAA (2017), and the corresponding DEM data was obtained from USGS data downloader USGS (2023). The aerial imagery from both events are 0.3 m $\times$ 0.3 m in resolution, the DEM data for Matthew 2016 are 6 m $\times$ 6 m and for Harvey 2017 are 3 m $\times$ 3 m in resolution. We resampled them into a resolution of 2 m $\times$ 2 m, which is fine enough for the purpose of flood extent mapping. Table 2 shows the regions we used, where R1–R5 are from Matthew 2016, and R6–R7 is from Harvey 2017. While high-resolution Earth imagery and DEM data are abundant, the ground-truth flood maps need to be annotated by domain experts. So, we provide a 3D visualization tool to domain experts for annotating pixels.
In our online appendix Sami et al. (2024), Figure 6 shows the imagery for Kinston, NC during Hurricane Matthew in 2016, its annotation where pixels are labeled as flooded or dry, and the normal-time imagery from Google Earth. We expect that inputting the normal-time image into a model is effective in helping identify flooded regions by providing a contrast. This is confirmed our experiments reported in Table 10 of our online appendix, which also aligns with findings from prior works Drakonakis et al. (2022); Rudner et al. (2019). Some pixels are ambiguous (e.g., covered with tree canopies) so remain unlabeled. Unlabeled pixels are not included in the loss computation during training, and not included during test when calculating the evaluation measures. As Table 2 shows, most pixels of each region are annotated. Without loss of generalization, in Table 2, we use R1 and R2 for training, and use R3, R4, R5, R6 and R7 for test.
@@
For physics-guided graphical models, we compare with hidden Markov tree (HMT) Xie et al. (2018); Jiang and Sainju (2019). Since HMT uses transductive learning, we use the same region for training and test. For U-Net and EvaNet, we use two regions R1 and R2 in Table 2 for training and the other regions for test. For example, when reporting the performance measures on R3, HMT is trained on R3 and tested on R3, while all other models are trained with the patches of R1 and R2, and then tested on patches of R3 and stitched.
By default, EvaNet uses only $\mathcal{L}_{eva}$ in Eq (2) as the loss function, where $w(\mathbf{p},\mathbf{p}_{n})$ uses the binary weighting scheme in Eq (4). We use 3 convolution blocks of Eva-Conv for both the encoder and the decoder as shown in Figure 4, where the input has 7 channels (disaster-time and normal-time RGB plus elevation map), and the encoder downsamples the spectral path using max pooling, and downsamples the elevation path using average pooling. These default configurations are found to be the best-performing, as demonstrated by our experiments in the online appendix Sami et al. (2024) (see Tables 610). We trained our EvaNet models and the U-Net baselines on a cluster with NVIDIA-P100 GPUs for 100 epochs with a learning rate of $1\times 10^{-7}$ and batch size of 4. These default settings are found to be the best-performing through extensive tests.
Performance Measures. Accuracy, precision, recall and F1-score are commonly used measures for binary classification. Following prior works such as Xie et al. (2018); Jiang and Sainju (2019), we consider our flood segmentation problem as a pixel-wise binary classification problem, and report the above four measures in two contexts: (1) “Dry” where dry (resp. flood) is the positive (resp. negative) class, and (2) “Flood” where dry (resp. flood) is the positive (resp. negative) class. These measures are computed only from labeled pixels in the test images.
Comparison with Baselines. Here, we consider 3 representative baselines: (1) HMT Xie et al. (2018), which is a graphical model for transductive learning that considers both the disaster-time RGB image and the elevation map, (2) U-Net 3C, which is a U-Net model that only considers the disaster-time RGB image, (3) U-Net 7C, which is a U-Net model that considers both the RGB images and the elevation map. They are compared with EvaNet 7C in Table 3. We skip 4C models since they are dominated by 7C ones (c.f. Table 10 in our online appendix Sami et al. (2024)).
@@
Ablation Study. To verify the effectiveness of both our Eva-Conv and Eva-Reg designs, we conduct an ablation study using 7-channel input, by exploring the performance of 4 models: (1) U-Net 7C (i.e., with loss $\mathcal{L}_{CE}$), (2) U-Net with our loss function $\mathcal{L}_{eva}$ (denoted by U-Net 7C + Eva-Reg), (3) U-Net 7C with our new Eva-Conv operations in both encoder and decoder (denoted by U-Net 7C + Eva-Conv), and (4) EvaNet 7C, i.e., with both Eva-Conv and Eva-Reg. Without loss of generality, Table 4 shows the performance results on test region R4, where we can see that (i) both Eva-Reg and Eva-Conv improve the model accuracy and F1-score as compared with U-Net 7C, (ii) Eva-Reg (i.e., our new loss $\mathcal{L}_{eva}$) improves the performance more significantly, and (iii) our EvaNet 7C model combining both designs provides the highest accuracy and F1-score. In our online appendix Sami et al. (2024), we show that our techniques are effective also in other ConvNet architectures using Table 11, and we also visually inspect the flood maps from different models in Figure 7 to confirm our conclusions.
Effect as a U-Net Drop-in Replacement. We also find that EvaNet can improve the performance of an existing model such as HMCT-PP Sainju et al. (2022) (i.e., U-Net + HMCT) by replacing its U-Net component. Due to space limit, please refer to Table 12 in our online appendix Sami et al. (2024) for more details.